word File Edit View Insert Format Tools Table Window Help EAN6 Fall2018 Home Insert Draw Design Layout References Mailings ReviewView Paste Pane WOEKSHeet 1. Finding the WACC You are given the following information concerning Happy Value Enterprise Common Stock: 225,000 shares of common stock selling for $64.50 per share. The stock has a beta of 0.85 and will pay a dividend of $2.70 next year. The dividend is expected to grow by 5 percent per year indefinitely. Preferred stock: 8,000 shares of 4.5 percent stock selling at $94 per share. Debt: a quoted price of 104. Market: tax rate. 9,000 62 percent coupon bonds outstanding, with 20 years to maturity and A 12 percent expected return, a 5 percent risk-free rate, and a 35 percent Calculate the WACC for Happy Value Enterprise. (7.53%) 2. Cost of Debt. Rinck Import has several bond issues outstanding, each making semiannual interest payments. The bonds are listed below. If the corporate tax rate is 34 %, what is the after-tax cost of the companys market-value debt? (4.69%) Bond Coupon rate Price Quote Maturity Face Value 7.5% 5.8 7.7 8.1 105.0 95.4 103.8 105.7 5 years 8 years 151/2 years 25 years $20,000,000 40,000,000 45,000,000 60,000,000 20 words English (United States)
Answer #1

Solution:-

1.

WACC of Happy value enterprises 10.24%

Explanation:-

For equity,

as per CAPM: Cost = Rf + (Rm - Rf) x Beta

  • Rf is the risk-free rate
  • Rm - Rf is the market premium of risk
  • Rm is the expected market return

Cost = 5.0 + (12.0 - 5.0) x 0.85 = 10.95%

For preferred stock:

  • Cost = D1 / P0 x 100.

Let face value be $100:

  • D1 = 4.6

Therefore, cost = 4.5 / 94.0 x 100 = 4.79%

Yield to maturity is the before-tax cost of debt.

Annual YTM = (C + (F - P) / n) / ((F + P ) / 2).

  • F = Face Value
  • P = Price
  • C = Periodic coupon
  • n = Number of periods

Let face value be $100:

  • (6.2 - 4.0 / 20) / (204.0 / 2) = 3.38%

Hence, annual YTM = 3.38 x 2 = 5.88%.

After-tax cost of debt = 5.88 x (1 - 0.35) = 3.822%

Total Market Value of Debt = 9000 x 104.0 = $936,000

Total Market Value of Preferred Stock = 8000 x 94.0 = $752,000

Total Market Value of Shares = 225000 x 64.50 = $14,512,500.00

Total Market Value of Capital Structure = $936,000.00 + $752,000.00 + $14,512,500.00 = $16,200,500.00

WACC:

  • 3.822 x 936,000 / 16,200,500 + 4.59 x 752,000 / 16,200,500 + 10.95 x 14,512,500/ 16,200,500 = 10.24%

Please Rate or comment if you have any doubt regarding this solution